B. (P-value Method). A teacher claims that more than 60% of the Grade 11 students preferred the use of the teacher-made video lessons as an alternative method in learning the Pre Calculus. In a survey of 200 students, 131 utilized the video lessons. Is there enough evidence to support the teacher's claim? Test the proportions at a-0.05.


1
Expert's answer
2021-07-06T14:30:44-0400

"P=0.6"

"x=131"

"n=200"

"\hat P=\frac{131}{200}=0.655"

"H_0:P=0.6"

"H_a:P>0.6"

"z=\frac{\hat P-P}{\sqrt{\frac{P(1-P)}{n}}}"

"=\frac{0.655-0.6}{\sqrt{\frac{0.6\times 0.4}{200}}}"

"1.5877"

"z_{0.05}=1.645"

Since 1.5877<1.645, we fail to reject the null hypothesis. There is no sufficient evidence to support the teacher's claim that more than 60% of the Grade 11 students preferred the use of the teacher-made video lessons as an alternative method in learning the Pre-Calculus.